# HouseLLM: LLM-Assisted Two-Phase Text-to-Floorplan Generation

### v3

@@ Abstract
This paper proposes a two-phase text-to-floorplan generation method, which guides a Large Language Model (LLM) to generate an initial layout (Layout-LLM) and refines them into the final floorplans through conditional diffusion model. We incorporate a Chain-of-Thought approach to prompt the LLM based on user text specifications, enabling a more user-friendly and intuitive house layout design. This method allows users to describe their needs in natural language, enhancing accessibility and providing clearer geometric constraints. The final floorplans generated by Layout-LLM through conditional diffusion refinement are more accurate and better meet user requirements. Experimental results demonstrate that our approach achieves state-of-the-art performance across all metrics, validating its effectiveness in practical home design applications. We plan to release our code for public use.
@@ 1 Introduction
Figure 1: Comparison of different floorplan generation paradigms. (a) Graph to house layout approach (e.g., HouseDiffusion), where rooms are represented as nodes and doors as edges, forming a graph that represents spatial relationships; (b) Text-to-floorplan approach, which directly maps a natural language description to a house layout; (c) Our two-phase paradigm, where an LLM is used to generate an initial layout, Layout-LLM, according to the user’s textual specification. The initial solution serves as a condition for generating the final layout, Layout-Final, through a diffusion model.
In architectural design, creating floorplans that meet user requirements remains a challenge. Traditional design methods rely not only on specialized knowledge but also require designers to make iterative adjustments to accommodate specific user needs, making personalized design difficult. Trained models [17, 15, 27, 13] have made efforts to improve the accuracy, interactivity, and efficiency of floorplans generation. Despite these advances, existing approaches have yet to achieve a level of user-friendliness and accuracy that enables widespread adoption by ordinary users.
The latest solutions [18, 19, 23] typically treat the floorplan creation task as a conditional generation problem. The conditions are expressed through a bubble diagram, where rooms are represented as nodes (or “bubbles”), and doors as edges specifying the spatial relationships between rooms, as shown in Figure 1(a). While the bubble diagram provides a precise way of describing the house layout, specifying a consistent graph structure [41, 39, 40] can still be too demanding for novice users who wish to explore designing with minimal effort. We believe that a more user-friendly interface should allow users to express their needs in natural language. For example, the user could simply specify “I need a house with three bedrooms, a living room, a bathroom, a kitchen, and a balcony adjacent to the living room.” Alternatively, the user may use a simple menu to make choices, which are then translated into natural language, perhaps in a more structured format, such as JSON. The text-to-floorplan paradigm would make the design process more accessible and intuitive for non-expert users.
A possible solution to the above requirement would be to use a generation model that directly maps text to house layouts, as illustrated in Figure 1(b). This end-to-end approach, however, faces significant challenges in meeting fine-grained geometric and numeric constraints specified in the text. Successfully training such models at least requires massive datasets, which are currently unavailable for floorplans. Indeed, our experiments will confirm the ineffectiveness of this approach even with the largest floorplan datasets available to date.
To address these issues, we propose a novel two-phase generation approach, called HouseLLM. In the first phase, we use a Multimodal Large Language Model (MLLM, or LLM for brevity) to interact with the user and translate the text to an initial design of house layout, termed Layout-LLM. In the second phase, we use a diffusion model to refine this initial design into a more accurate, final house layout, termed Layout-Final. This two-phase generation paradigm is depicted in Figure 1(c).
Drawing on the power of LLMs, our method provides a natural user interface and offers basic floorplan generation capabilities without the need for dedicated training. To leverage the LLM, we propose a prompting approach based on the Chain-of-Thought (CoT) technique [30, 38]. This approach enables generation of a house layout that meets the user’s core requirements regarding the number of rooms, room types, and approximate positions, although it may have imperfections in object sizing and alignment. To further optimize the result, we have designed a diffusion model that refines the initial layout into a more accurate and detailed final floorplan. The model basically follows the design of HouseDiffusion, with a key difference: while HouseDiffusion applies conditioning only in the reverse process, our approach incorporates conditioning in both the forward (noise addition) and reverse processes, allowing for more robust alignment with the specified conditions.
@@
We have validated our approach on the RPlan dataset, in comparison with state-of-the-art methods. For instance, in comparison with HouseDiffusion [23], our method achieves the best performance across all metrics, and in particular, with a 28% improvement in diversity. To summarize, this paper makes the following contributions:
  1. 1.
    We present a novel two-phase floorplan generation approach that leverages LLM to greatly reduce the burden of learning from scratch as adopted by previous work.
  2. 2.
@@
    We propose prompt designs based on the chain-of-thought technique. The prompts successfully guide the LLM to generate initial house layouts.
  3. 3.
    We design a conditional diffusion model to refine the layouts generated by the LLM.
  4. 4.
    The experimental results show that our method achieves state-of-the-art performance.
@@ 2 Related Work
Figure 2: Training and testing processes of our method. (a) An example of LLM generating a Layout-LLM according to user demands. (b) Given a house layout sample, we use the LLM to describe it. The description is used to mimic the user’s demands. Using multiple examples as in (a) as demos, we ask the LLM to generate a Layout-LMM for each sample. These initial layouts serve as conditions of the generator, which outputs Layouts-Finals. (c) Given a textual description from the user, we again use the demos like (a) to obtain a Layout-LLM, which goes through the diffusion model to generate Layout-Final.
@@
Floorplan generation. In the field of building design, generating high-quality house layouts has been an important research direction [5, 33, 10, 16, 20, 27], and many innovative approaches have emerged in recent years. Nauata et al. [18] proposed House-GAN, a method based on Generative Adversarial Networks (GANs) to generate house layouts. This method achieves end-to-end automated house layout generation. Nauata et al. [19] further proposed House-GAN++, which improved the original GAN structure with enhanced generation quality. In particular, House-GAN++ addressed the problems of missing doors, making the generated houses more practical. Upadhyay et al. [28] expanded on this by considering user inputs in the form of boundary, room types, and spatial relationships. Hu et al. [10] presented Graph2Plan, which retrieves a set of floorplans with their associated layout graphs from a database, allowing a user to specify room counts and other layout constraints. Shabani et al. [23] proposed the HouseDiffusion model, which introduced bubble diagrams as constraints and used a diffusion model to generate house layouts. Recently, Chen et al. [3] transformed visual sensor data into polygonal shapes with Diffusion Models. Su et al. [26] learn topology graphs via its bi-directional structure of ‘corruption and denoise’ is developed.
@@
Conditional Diffusion. Conditional diffusion models are a subset of diffusion models [2, 24, 34] where the generation is conditioned on specific input data, such as labels, images, or text, allowing more control over the output [37, 4, 8]. Expanding on the success in diffusion model [24], Ho et al. proposed DDPMs [25], introducing constraints within the diffusion process to guide generation, which led to significant improvements success. Subsequently, conditional diffusion models DALLE-2 [21] and Stable Diffusion [22] have been applied to text-to-image generation. Recent advancements continue to refine diffusion models’ conditional control. For instance, Yang et al. [35] improved diffuse image synthesis based on context prediction. Yang et al. [36] proposed an end-to-end lossy image compression framework using conditional diffusion models to refine the source image. The conditional diffusion model is a foundation of our approach, in which use it to refine the floorplan generated by the LLM.
@@ 3 Method
HouseLLM uses a two-phase approach to generate house layouts. By breaking the generation pipeline into two phases, HouseLLM manages to exploit the power of LLMs in interpreting user demands and generating approximate layouts with their common knowledge and reasoning ability. Further refinement is designed to enhance the layout’s quality. Figure 2 depicts the training and testing processes of our method, in which Layout-LLM plays a pivotal role.
3.1 Overall framework and Layout-LLM
Figure 3: CoT-based prompting for generation of Layout-LLM, and its function in training and testing. (a) An example showing how to interact with the LLM to obtain Layout-LLM. The Initialization section defines the LLM’s role as a house designer, and standardizes output format; the Chain of Thought section directs the LLM to create a house layout step by step, ensuring plausible room placement and sizing. (b) Invoking the Layout-LLM generation during training and testing.
Natural language descriptions alone may lack the detailed context necessary for accurate reasoning about specific house layouts [14, 31]. To address this limitation, we enhance the LLM’s reasoning capability by allowing it to reference multiple human-designed layout demos, following the Chain of Thought (CoT) approach. By demonstrating each step explicitly, including room selection, placement, and sizing decisions, CoT helps the LLM break down complex layout generation tasks into manageable, step-by-step inferences.
The demoset $E$ for the prompt engineering consists of pairs of language description $e_{d}$ and corresponding Layout-LLM $e_{l}$:
$E=\{(e_{l},e_{d})_{1},(e_{l},e_{d})_{2},...,(e_{l},e_{d})_{n}\}.$ (1)
@@ 3.1 Overall framework and Layout-LLM
$\text{House}=\{\text{rooms}=[\text{Room}_{\text{1}},...,\text{Room}_{\text{n}}]\}\vspace{-5pt}$ (2)
@@
$\text{door}=[\text{Direction}_{\text{i}}]\}$
@@
The training process requires establishing a mapping from ground-truth layouts to Layout-LLMs, which are further used as conditions for generating Layout-Finals. Given a house layout sample from the dataset, we first use the LLM to give a description of the layout, which mimics the user’s demands. The description is then fed to the LLM, which generates Layout-LLMs using the CoT-based prompt.
Figure 3 shows a sample of prompts design. The prompt used in our method consists of three parts: Initialization, Chain of Thought (CoT), and Layout-LLM generation.
3.2 Generating Layout-Final
Figure 4: Conditional diffusion network for refining Layout-LLM. The forward process takes the ground-truth house layout $x^{g}_{0}$ and the Layout-LLM $x^{l}_{0}$ and adds a Gaussian noise to create a noisy house layout sample $x_{T}$. The reverse process takes a noisy house layout at time $t$ with Layout-LLM as the condition. Two encoders are used to encode and obtain the latent representations for $x^{g}_{t}$ and $x^{l}_{t}$.
Following HouseDiffusion [23], our conditional generation method is inspired by the diffusion model. The reverse process consists of a set of modules: Layout-LLM Encoder, Noisy Encoder, Continuous Decoder and Discrete Decoder. Figure 4 presents the specific training method of our model.
Unlike HouseDiffusion [23], which applies conditional information only during noise decoding, our approach fuses conditional information in both noise addition and decoding phases. By embedding Layout-LLM from the start, the model builds condition-relevant features progressively during noise accumulation and continues to refine them in the denoising phase. This dual-phase conditioning enhances condition alignment and improves the stability and accuracy of the generated layouts.
@@ 3.2.1 Diffusion Model
Diffusion models progressively refine Gaussian noise $x_{T}$ over $T$, guided by a conditional Layout-LLM $y$. The forward process takes an initial data sample $x_{0}$, a conditional Layout-LLM $y$ and generates a noisy sample $x_{t}$ at time step $t$ by sampling Gaussian noise $\epsilon\sim N(0,1)$, incorporating information from the conditional layout through its embedding $e(y)$:
$x_{t}=\sqrt{\alpha_{t}}*x_{0}+\sqrt{1-\alpha_{t}}*\epsilon+e(y)$ (4)
$\alpha_{t}$ represents the noise schedule, controlling noise intensity changes from $1$ to 00 at each time step $t$. $e(y)$ is the embedding derived from the Layout-LLM $y$, guiding the diffusion model to align with the features of the Layout-LLM $y$.
@@
$\hat{x}_{0}=f_{\theta}(x_{t},t,e(y))$ (5)
$f_{\theta}$ is the denoising network Transformer [29] parameterized by $\theta$. $x_{t}$ is the noisy image at current time step $t$.
The reverse process updates the state of $x_{t}$ each step to remove noise and gradually guide the generated image toward the structure of the condition $y$.
$x_{t-1}=\sqrt{\alpha_{t-1}}*\hat{x}_{0}+\sqrt{1-\alpha_{t-1}}*\epsilon$ (6)
@@
3.2.2 Network
Data representation. Here we refer to the processing of data representation in [23] to meet the constraints on the number and type of rooms. Suppose $P^{g}=\{P^{g}_{1},P^{g}_{2},...,P^{g}_{N}\}$ is the set of polygonal loops for each room/door in the original data set. $P^{l}=\{P^{l}_{1},P^{l}_{2},...,P^{l}_{N}\}$ is for Layout-LLM, and $P^{f}=\{P^{f}_{1},P^{f}_{2},...,P^{f}_{N}\}$ is for the LLM-Final. For $\forall P\in\{P^{g},P^{l},P^{f}\}$, each $p_{i}$ is defined by a sequence of corners with as a 2-dim coordinates:
$P_{i}=\{C_{i,1},C_{i,2},...,C_{i,N_{i}}|C_{i,j}\in R^{2}\}$ (7)
Latent Encoder. Given a floorplan sample $P^{g}_{t}$ and corresponding Layout-LLM $P^{l}_{t}$ at time $t$, every room/door corner $C_{i,j}^{t}$ is mapped to $\hat{C}^{t}_{i,j}$ through the Layout-LLM and the Noisy encoder:
$\mbox{Linear}([\mbox{AU}(C^{g}_{i,j})_{t},\mbox{AU}(C^{l}_{i,j})_{t},R_{i},\mathbf{1}(i),\mathbf{1}(j),t])$ (8)
Following the method in [23], AU enhances relational inference by inserting angular coordinates. $R_{i}$ represents a 25-dimensional one-hot vector indicating room type. $1(\centerdot)$ refers to a 32-dimensional one-hot vector for the room index $i$ and a corner index $j$. The scalar $t$ is included, and a linear layer transforms the embedding into a 512-dimensional vector.

### v4

# HouseTune: Two-Stage Floorplan Generation with LLM Assistance

@@ Abstract
This paper proposes a two-stage text-to-floorplan generation framework that combines the reasoning capability of Large Language Models (LLMs) with the generative power of diffusion models. In the first stage, we leverage a Chain-of-Thought (CoT) prompting strategy to guide an LLM in generating an initial layout (Layout-Init) from natural language descriptions, which ensures a user-friendly and intuitive design process. However, Layout-Init may lack precise geometric alignment and fine-grained structural details. To address this, the second stage employs a conditional diffusion model to refine Layout-Init into a final floorplan (Layout-Final) that better adheres to physical constraints and user requirements. Unlike prior methods, our approach effectively reduces the difficulty of floorplan generation learning without the need for extensive domain-specific training data. Experimental results demonstrate that our approach achieves state-of-the-art performance across all metrics, which validates its effectiveness in practical home design applications. Our code will be made publicly available.
In architectural design, creating floorplans that align with user requirements remains a challenge. Traditional design methods not only rely on specialized knowledge but also require designers to make iterative adjustments to accommodate specific user needs, which makes personalized design challenging. Learning-based models Murali et al. (2017); Merrell et al. (2010); Sun et al. (2022); Luo and Huang (2022) have made efforts to improve the accuracy, interactivity, and efficiency of floorplan generation. Despite these advances, existing approaches have yet to achieve a level of user-friendliness and accuracy that enables ready adoption by ordinary users.
Existing solutions Nauata et al. (2020, 2021); Shabani et al. (2023) typically treat the floorplan creation task as a conditional generation problem. The conditions are expressed through a bubble diagram, where rooms are represented as nodes (or “bubbles”), and doors as edges specifying the spatial relationships between rooms, as shown in Figure 1(a). While the bubble diagram provides a precise way of describing the house layout, specifying a consistent graph structure can still be too demanding for novice users who wish to explore designing with minimal effort. We believe that a more user-friendly interface should allow users to express their needs in natural language. For example, the user could simply specify “I need a house with three bedrooms, a living room, a bathroom, a kitchen, and a balcony adjacent to the living room.” Alternatively, the user may use a simple menu to make choices, which are then translated into natural language. The text-to-floorplan paradigm make the design process more accessible and intuitive for non-expert users.
A solution to this requirement involves training a generative model capable of directly mapping textual descriptions to house layouts, as depicted in Figure 1(b). The Tell2Design method Leng et al. (2023) addresses this task using a Sequence-to-Sequence approach, where the input text specifies the floorplan boundary and the exact geometry of individual rooms. The method imposes stringent demands for geometric detail and consistency, placing an even greater burden on users compared to working with bubble diagrams. In addition, the design limits the diversity of the results, which is often crucial for exploratory design.
@@ 1 Introduction
Figure 1: Comparison of different floorplan generation pipelines. (a) Graph-to-floorplan approach (e.g., HouseDiffusion), where rooms are represented as nodes and doors as edges, forming a graph Zhuo and Tan (2022a, b); Zhuo et al. (2024) that represents spatial relationships; (b) Text-to-floorplan approach, which directly maps a natural language description to a house layout; (c) Our two-stage pipeline, where an LLM is used to generate an initial layout, Layout-Init, according to the user’s textual specification. The initial solution serves as a condition for generating the final layout, Layou-Final, through a diffusion model.
We introduce HouseTune, a novel two-stage floorplan generation framework. In the first stage, we use a multimodal large language model (LLM) to generate an initial house layout, termed Layout-Init. The LLM’s common-sense knowledge and reasoning capability enable the creation of a preliminary floorplan based on user input. In the second stage, a diffusion model refines this initial design into a more precise and reasonable final layout, referred to as Layout-Final. This two-stage generation process is illustrated in Figure 1(c).
Generating a reasonable house layout using an LLM is not straightforward, as LLMs often struggle to interpret exact numeric and geometric constraints. To address this, we employ a Chain-of-Thought (CoT) prompting strategy Wei et al. (2022); Zhang et al. (2022). This approach ensures the generated layout aligns with the user’s core requirements, such as the number of rooms, room types, and approximate spatial arrangements.
The initial layouts produced at this stage typically exhibit imperfections in object sizing and alignment, due to the inherent limitations of LLMs in handling intricate geometric details. To overcome this, we design a diffusion model that refines the initial layout into a final floorplan. While inspired by HouseDiffusion Shabani et al. (2023), our approach differs in its conditioning strategy: whereas HouseDiffusion applies conditioning solely during the reverse process, our model integrates conditioning in both the forward (noise addition) and reverse processes. This dual-conditioning mechanism ensures enhanced alignment with the specified constraints, yielding more precise and robust results.
@@
We validate our approach on the RPlan dataset, in comparison with state-of-the-art methods. For instance, in comparison with HouseDiffusion, our method achieves the best performance across all metrics, and in particular, with a 28% improvement in diversity. To summarize, this paper makes the following contributions:
  • We propose a novel two-stage floorplan generation framework that leverages the initial layouts generated by LLMs as the condition to generate the final floorplans.
@@
  • We develop a prompt design based on the Chain-of-Thought technique, successfully guiding the LLM to produce structured and coherent initial house layouts.
  • We design a conditional diffusion model that refines the LLM-generated layouts into high-fidelity floorplans. Experimental results demonstrate that our method achieves state-of-the-art performance.
@@ 2 Related Work
Figure 2: Training and testing processes of our method. (a) An example of LLM generating a Layout-Init according to user demands. (b) Given a house layout sample, we use the LLM to describe it. The description is used to mimic the user’s demands. Using multiple examples as in (a) as demos, we ask the LLM to generate a Layout-Init for each sample. These initial layouts serve as conditions for the generator, which outputs Layouts-Final. (c) Given a textual description from the user, we again use the demos like (a) to obtain a Layout-Init, which goes through the diffusion model to generate Layout-Final.
@@
Floorplan Generation. In the field of building design, generating high-quality house layouts has been an important research direction Hendrikx et al. (2013); Wu et al. (2018); Hu et al. (2020); Müller et al. (2006); Peng et al. (2014); Sun et al. (2022). Nauata et al. Nauata et al. (2020) proposed House-GAN, a method based on Generative Adversarial Networks (GANs) that achieves end-to-end automated house layout generation. Nauata et al. Nauata et al. (2021) further proposed House-GAN++, which improved the original GAN structure by addressing the problems of missing doors. Upadhyay et al. Upadhyay et al. (2022) expanded on this by considering user inputs in the form of boundaries, room types, and spatial relationships. Hu et al. Hu et al. (2020) presented Graph2Plan, which retrieves a set of floorplans with their associated layout graphs from a database, allowing a user to specify room counts and other layout constraints. Shabani et al. Shabani et al. (2023) introduced bubble diagrams as constraints and used a diffusion model to generate house layouts. Chen et al. Chen et al. (2024) transformed visual sensor data into polygonal shapes with Diffusion Models. Su et al. Su et al. (2024) developed a bi-directional structure of “corruption and denoise” approach to learn topology graphs. Zeng et al. Zeng et al. (2024) proposed a multi-conditional two-stage generation model, which allows human designers to intervene and enhance controllability based on the denoising diffusion model. Leng et al. Leng et al. (2023) introduced the Tell2Design (T2D) dataset, which comprises over 80,000 floorplan designs paired with natural language descriptions, in support of floorplan generation. HOLODECK Yang et al. (2024a) explored the use of LLMs to generate floorplans populated with various objects, emphasizing the overall quality and consistency of the environments rather than dealing with complex floorplans.
@@
Conditional Diffusion. Conditional diffusion models are a subset of diffusion models Cao et al. (2024); Sohl-Dickstein et al. (2015); Yang et al. (2023) where the generation is conditioned on specific input data, such as labels, images, or text, allowing more control over the output Zhang et al. (2023); Chen et al. (2022); Ho et al. (2022). Expanding on the success of diffusion models Sohl-Dickstein et al. (2015), Ho et al. proposed DDPMs Song et al. (2020), introducing constraints within the diffusion process to guide generation, which led to significant significant improvements. Subsequently, conditional diffusion models, such as DALLE-2 Ramesh et al. (2022) and Stable Diffusion Rombach et al. (2022) have been applied to text-to-image generation. Recent advancements continue to refine diffusion models’ conditional control. For instance, Yang et al. Yang et al. (2024b) improved diffuse image synthesis based on context prediction. Yang et al. Yang and Mandt (2024) proposed an end-to-end lossy image compression framework using conditional diffusion models to refine the source image. The conditional diffusion model is a foundation of our approach, in which use it to refine the floorplan generated by the LLM.
@@ 3 Method
HouseTune uses a two-stage approach to generate house layouts. By breaking the generation pipeline into two stages, HouseTune manages to exploit the power of LLMs in interpreting user demands and generating approximate layouts with their common knowledge and reasoning ability. Further refinement is designed to enhance the layout’s quality. Figure 2 depicts the training and testing processes of our method, in which Layout-Init plays a pivotal role.
3.1 Layout-Init Generation
Figure 3: CoT-based prompting facilitates generation of Layout-Init and its function in training and testing. (a) An example showing how to interact with the LLM to obtain Layout-Init. The Initialization section defines the LLM’s role as a house designer, and standardizes output format; the Chain of Thought section directs the LLM to create a house layout step by step, ensuring plausible room placement and sizing. (b) Invoking the Layout-Init generation during training and testing.
Natural language descriptions alone may lack the detailed context necessary for accurate reasoning about specific house layouts Mann et al. (2020); Work . To address this limitation, we enhance the LLM’s reasoning capability by allowing it to reference multiple human-designed layout demos, following the Chain of Thought (CoT) approach. By demonstrating each step explicitly, including room selection, placement, and sizing decisions, CoT helps the LLM break down complex layout generation tasks into manageable, step-by-step inferences.
The demo set $E$ for the prompt engineering consists of pairs of language description $e_{d}$ and corresponding Layout-Init $e_{i}$:
$E=\{(e_{i},e_{d})_{1},(e_{i},e_{d})_{2},...,(e_{i},e_{d})_{n}\}.$ (1)
@@ 3.1 Layout-Init Generation
$\text{House}=\{\text{rooms}=[\text{Room}_{\text{1}},...,\text{Room}_{\text{n}}]\},\vspace{-5pt}$ (2)
@@
$\text{door}=[\text{Direction}_{\text{i}}]\},$
@@
The training process requires establishing a mapping from ground-truth layouts to Layout-Init, which are further used as conditions for generating Layout-Final. Given a house layout sample from the dataset, we first use the LLM to give a description of the layout, which mimics the user’s demands. The description is then fed to the LLM, which generates Layout-Init using the CoT-based prompt.
Figure 3 shows a sample of prompt design. The prompt used in our method consists of three parts: Initialization, Chain of Thought (CoT), and Layout-Init generation.
Initialization. Initialization introduces task settings, input and output format specifications, and reasoning sub-tasks. Specifically, the task utilizes the role-prompting technique John (2023) to instruct the LLM to play the role of a designer: we present specific tasks to the LLM, prompting it to explain, step by step, the reasoning behind the design of each house location. An example prompt illustrating this setup is provided in see Figure 3 (a).
Chain of thought. In each separate reasoning step, the LLM needs to determine the current house layout, then randomly select the next room to be generated, and infer the information for the next room based on the information from the existing rooms. The emphasis is on explaining the rationale behind these decisions. Figure 3 (a) presents an example that shows how to guide the model through each step of the inference process, including room selection, sizing, and positioning within the layout.
Layout-Init generation. In this step, demos are input into the LLM to prompt the generation of the house layout. The user provides natural language input, instructs the LLM to perform a step-by-step reasoning process. The final output is structured in the JSON format.
3.2 Diffusion Refinement for Layout-Final
Figure 4: Conditional diffusion network for refining Layout-Init. The forward process takes the ground-truth house layout $x^{g}_{0}$ and the Layout-Init $x^{i}_{0}$ and adds a Gaussian noise to create a noisy house layout sample $x_{T}$. The reverse process takes a noisy house layout at time $t$ with Layout-Init as the condition. Two encoders are used to encode and obtain the latent representations for $x^{g}_{t}$ and $x^{i}_{t}$.
Following HouseDiffusion Shabani et al. (2023), our conditional generation method is inspired by the diffusion model. The reverse process consists of a set of modules: Layout-Init Encoder, Noisy Encoder, Continuous Decoder and Discrete Decoder. Figure 4 presents the specific training method of our model.
Unlike HouseDiffusion Shabani et al. (2023), which applies conditional information only during noise decoding, our approach fuses conditional information in both noise addition and decoding phases. By embedding Layout-Init from the start, the model builds condition-relevant features progressively during noise accumulation and continues to refine them in the denoising phase. This dual-phase conditioning enhances condition alignment and improves the stability and accuracy of the generated layouts.
@@ 3.2.1 Diffusion Model
Diffusion models progressively refine Gaussian noise $x_{T}$ over $T$, guided by a conditional Layout-Init $y$. The forward process takes an initial data sample $x_{0}$, a conditional Layout-Init $y$ and generates a noisy sample $x_{t}$ at time step $t$ by sampling Gaussian noise $\epsilon\sim N(0,1)$, incorporating information from the conditional layout through its embedding $e(y)$:
$x_{t}=\sqrt{\alpha_{t}}*x_{0}+\sqrt{1-\alpha_{t}}*\epsilon+e(y).$ (4)
$\alpha_{t}$ represents the noise schedule, controlling noise intensity changes from $1$ to 00 at each time step $t$. $e(y)$ is the embedding derived from the Layout-Init $y$, guiding the diffusion model to align with the features of the Layout-Init $y$.
@@
$\hat{x}_{0}=f_{\theta}(x_{t},t,e(y)).$ (5)
$f_{\theta}$ is the denoising network Transformer Vaswani (2017) parameterized by $\theta$. $x_{t}$ is the noisy image at current time step $t$.
The reverse process updates the state of $x_{t}$ each step to remove noise and gradually guide the generated image toward the structure of the condition $y$:
$x_{t-1}=\sqrt{\alpha_{t-1}}*\hat{x}_{0}+\sqrt{1-\alpha_{t-1}}*\epsilon.$ (6)
@@
3.2.2 Network Architecture
Data representation. Here we refer to the processing of data representation in Shabani et al. (2023) to meet the constraints on the number and type of rooms. Suppose $P^{G}$ is the set of polygonal loops for each room/door in the ground truth. $P^{I}$ is for Layout-Init, and $P^{F}$ is for the LLM-Final. For $\forall P\in\{P^{G},P^{I},P^{F}\}$, each loop $P_{i}$ is defined by a sequence of corners with as a 2-dim coordinates:
$P_{i}=\{C_{i,1},C_{i,2},...,C_{i,N_{i}}|C_{i,j}\in R^{2}\}.$ (7)
Latent Encoder. Given a floorplan sample $P^{G}_{t}$ and corresponding Layout-Init $P^{I}_{t}$ at time $t$, every room/door corner $C_{i,j}^{t}$ is mapped to $\hat{C}^{t}_{i,j}$ through the Layout-Init and the Noisy encoder:
$\mbox{Linear}([\mbox{AU}(C^{t}_{i,j})_{G},\mbox{AU}(C^{t}_{i,j})_{I},R_{i},\mathbf{1}(i),\mathbf{1}(j),t]).$ (8)
Following the method in Shabani et al. (2023), AU enhances relational inference by inserting angular coordinates. $R_{i}$ represents a 25-dimensional one-hot vector indicating room type. $1(\cdot)$ refers to a 32-dimensional one-hot vector for the room index $i$ and a corner index $j$. The scalar $t$ is included, and a linear layer transforms the embedding into a 512-dimensional vector.